the area of a parallelogramis 72cm2. if one of its side is 8cm and one of its height is 6cm, find the other height and perimeter. if managed to solve for the area but cannnot solve the perimeter. i am having troubles solving questions requiring students to find the perimeter but the height or side is not totally provided
Answer by DrBeeee(684)

Use the formula for the area of a parallelogram twice.
(1) A = bh
The first time we know the area, A, is 72, and the height, h, is 6, giving us
(2) 72 = b*6 or
(3) b = 12, which is the lenght of the other two sides.
The second time we know the area, A, is 72, and the base, b, is 8, giving us
(4) 72 = 8*h or
(5) h = 9, which is the other height, or perpendicular distance between the parallel sides of length 8.
Now you have the lenght of all four sides, just use the formula for the perimeter,
(6) p = sum of four sides or
(7) p = 8 + 8 + 12 + 12 or
(8) p = 40
Answer: The given parallelogram has heights of 6cm and 9cm and its perimeter is 40cm.